OpenStudy (anonymous):

Help making this--> r=3sin2theta<-- into x and y to graph? I'm stuck.

OpenStudy (anonymous):

@whpalmer4 ?

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

\[x=r \cos \theta, y=r \sin \theta ,\] \[x^2+y^2=r^2\left( \cos ^2\theta+\sin ^2 \theta \right)=r^2\] \[r=3 \sin 2 \theta =3*2 \sin \theta \cos \theta \] \[r^3=6 r \sin \theta * r \cos \theta\] \[\left( \sqrt{x^2+y^2} \right)^3=6 y x\] \[\left( x^2+y^2 \right)^3=36x^2y^2\]

OpenStudy (whpalmer4):

Point of clarification: is it \[r = 3\sin 2\theta\]or\[r = 3\sin^2\theta\]?

OpenStudy (anonymous):

if \[r=3 \sin ^2\theta =3*\frac{ y^2 }{ r^2 },r^3=3 y^2,\left( \sqrt{x^2+y^2} \right)^3=3y^2\] \[\left( x^2+y^2 \right)^3=9y^4\]

OpenStudy (anonymous):

No, the 2 was big beside theta, but the answer choices are 4 graphs.. how do I plug what you gave me into a graph?

OpenStudy (anonymous):

@phi ?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@sourwing ?

OpenStudy (phi):

to find the graph (by hand) is make a table angle 2*angle r=3 sin(2A) 0 0 0 15 30 1.5 22.5 45 2.1 30 60 2.6 45 90 3 60 120 2.6 67.5 135 2.1 75 150 1.5 90 180 0 you should keep going for the angles in the other quadrants (you should get the same pattern) to plot these points, draw a line at each angle. make each line as long a "r" draw a smooth curve connecting the tips of the lines